OpenStudy (anonymous):

find the 4th term of the expansion of (x-4y)^7

OpenStudy (raden):

it is 7C3 x^(7-3)(-4y)^3 then simplify

OpenStudy (kenljw):

1 1 1 1 2 1 2 1 3 1 3 3 1 4 1 4 6 4 1 5 1 5 10 10 5 1 6 1 6 15 20 15 6 1 7 (x+(-4y))^7 forth term 20 x^(7-3) (-4y)^3 = -20X4^3 x^4 y^3 The triangle I derived is easy to develop it's called the Pascal Triangle and is used for coefficients of a binomial.

OpenStudy (raden):

7C3 = 7!/(7-3)!3! = 7!/4!3! = 7.6.5.4!/4!3! = 7 * 5 = 35

OpenStudy (anonymous):

then wat about (x-2y)^4, the 4th term in expansion?

OpenStudy (kenljw):

(x+(-2y))^4 there are four terms (-2y)^4 = 16y^4

OpenStudy (kenljw):

sorry 5 terms x(-2y)^3 = -8xy^3

OpenStudy (kenljw):

for 7 and need to add one more line to trangle 1 7 21 35 35 21 7 1 35 x^(7-3) (-2)y^3 = 35X(-2)^3x^4y^3